Two Block of BCC iron 1 mole each are kept in thermal contact with each other and they are packed in an adiabatic enclosure first block wast at 300 k and second was at 900 k to start with what should be entropy produced ?


1
Expert's answer
2021-02-22T05:53:37-0500

Tf = ½ (T1+T2) = ½ (300+900) = 600 K

ΔS1 = nCp ln (Tf / T1) = Cp ln (Tf / T1) – 1 mol

ΔS2 = nCp ln (Tf / T2) = Cp ln (Tf / T2) – 1 mol

ΔS total = ΔS1 + ΔS2 = Cp ln (Tf2 /T1T2)

Cp = 24.3 J/mol K

ΔS total = 24.3 ln(6002/300x900) = 6.99